# Data-Driven Characterization of Latent Dynamics on Quantum Testbeds

### v1

@@ 4.2 Learning Dynamics of QPU - Dev2
Having demonstrated the effectiveness of the UDE approach on a more stable Dev1 QPU, we investigate its predictive accuracy when applied to the more noisy Dev2 QPU. In this analysis, we explore the effectiveness of both an Experiment-Generalized operator, which models the dynamics across all experiments in expectation, and an Experiment-Specific operator designed for a single experiment.
We first apply our procedure to learn Exp-Gen UDE models using training data over a $10\mu s$ window from five experiments (i.e. $|\mathrm{\boldsymbol{\xi}}|=5$ in Eq. 12). Figure 6 shows the probability distribution of the trace distance, and Table 4 summarizes the corresponding statistical moments for each model. The distribution of the trace distance shows that the learned UDE models outperform the base model in accuracy over both the training and validation data sets. Here we observe that the affine and the nonlinear neural network models exhibit superior performance when compared to the structure preserving ansatz. This suggests that the Markovian Lindblad model is unable to accurately model all processes present on the QPU, and these processes are better represented by the affine and nonlinear operators. By comparing Tab. 2 and Tab. 4, it is apparent that the learned UDE models capture dynamics on both QPUs to similar accuracy in expectation, but variance on Dev2 is an order of magnitude higher than it is on Dev1. This is attributed to the increased levels of noise present on this QPU. Nevertheless, this demonstrates that our UDE approach is capable of effectively capturing the dynamics of noisy QPUs.
@@
where the magnitude of detuning is similar to those obtained on Dev1. The magnitude of the perturbation for $\sigma_{x}$ and $\sigma_{y}$ is, 1.96 kHz and 10.4 kHz respectively, and are significantly larger than on Dev1. The estimated perturbations to the decoherence times are $\mathcal{T}^{Lind}=\left\{10,8.5\right\}~\mu s$. The perturbed $T_{1}$ decay time is approximately $4.6\mu s$ and is an order of magnitude lower than the decay time estimated using standard characterization protocols. The effective $T_{2}$ dephasing time is approximately $1.6\mu s$ and also lower than what was obtained through standard characterization.
Having demonstrated the UDE approach’s ability to learn generalized operators, we investigate the accuracy gained by tailoring the operator for a particular experiment. Figure 7 displays the expected energy, and probability density of the trace distance for a random experiment; Tab. 5 presents the statistical moments for the different UDE approaches when learning Experiment-Specific operators. We see that the expected energy of the density matrices computed by the affine and nonlinear models is more closely aligned with the those obtained from LIE; furthermore, the probability densities are higher and more concentrated for smaller trace distances. Although the accuracy of the Exp-Spec operators over the training data is greater than that of the Exp-Gen operators, the accuracy of the two over the validation set are comparable. The expected energy and densities of the trace distance for additional random experiments can be found in Appendix A; these results demonstrate improved accuracy. This reinforces the conclusion that learning Experiment-Specific operators greatly improves accuracy.
@@
where the detuning introduced by the Exp-Spec operator is significantly lower than that estimated by the Exp-Gen operator. The estimated perturbed decoherence times are $\mathcal{T}^{Lind}=\left\{17.9,23.1\right\}~\mu s$, with the perturbed $T_{1}$ decay and $T_{2}$ dephasing times of approximately $7.8\mu s$ and $2.9\mu s$ respectively. The lower $T_{1}$ estimates agree with the fast decaying nature of the expected energy seen in Fig. 7, where for $N=2$, the expected energy is equivalent to the population of the first excited state.
@@ 5 Conclusion
This work demonstrated a data-driven approach for learning latent dynamics of quantum processing units (QPUs). The technique augments a base model of the known dynamics (e.g. Lindblad master equation) with a trained source term that models the unknown dynamics apparent in the training data from QPUs. We presented analysis with two base models: Liouville-von Neumann and Lindblad master equation, and three formulations (ansatze) for this data-driven source term: a structure preserving model, an affine model and a nonlinear neural network model. The structure preserving ansatz allowed for both model interpretability and offered physically consistent time-evolution of the quantum state, whereas the affine and nonlinear models were not directly interpretable nor yielded physically consistent evolution. The affine and nonlinear models were made physically consistent by applying a spectral filter with renormalization, a-posteriori. The accuracy and applicability of the different ansatze was investigated using data from two different QPUs at the LLNL’s Quantum Design and Integration Testbed (QuDIT), each with a different level of noise.
We demonstrated that the structure preserving and affine models typically required less training data than the nonlinear model while achieving improved out-of-distribution accuracy over the base models. Furthermore, the structure preserving model converged to similar operators regardless of the underlying base model used. We demonstrated that the structure preserving ansatz is more accurate for less-noisy QPUs where the latent dynamics are well described by the Markovian Lindblad equation, whereas the affine model performed well on both QPUs. The nonlinear model was seen to perform well only on more noisy QPUs, and hence, both the affine and nonlinear models are better able to model the noise processes. Despite the lack of interpretability and for general application of learning dynamics and noise processes, the affine model was seen to be more accurate and robust, than the remaining ansatze. Both structure preserving and affine models are applicable for learning dynamics of stable, less-noisy QPUs, and can be used as more accurate models for solving optimal control problems.
@@ Acknowledgements
This work was performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under Contract DE-AC52-07NA27344. LLNL-JRNL-858483.
@@ Appendix A Supplementary Material: Learning Dynamics of QPU - Dev2
and the estimated perturbed decoherence times are $\mathcal{T}=\left\{11.2,11.1\right\}~\mu s$ with the perturbed $T_{1}$ decay and $T_{2}$ dephasing times of approximately $5.1\mu s$ and $1.9\mu s$ respectively.
@@
and the estimated perturbed decoherence times are $\mathcal{T}=\left\{7.8,6.2\right\}~\mu s$ with the perturbed $T_{1}$ decay and $T_{2}$ dephasing times of approximately $3.7\mu s$ and $1.2\mu s$ respectively.

### v2

@@ 4.2 Learning Dynamics of QPU - Dev2
Having demonstrated the effectiveness of the UDE approach on a more stable Dev1 QPU, we investigate its predictive accuracy when applied to the more noisy Dev2 QPU. In this analysis, we explore the effectiveness of both an Experiment-Generalized and Experiment-Specific operators.
We first apply our procedure to learn Exp-Gen UDE models using training data over a $10\upmu$s window from five experiments (i.e. $|\mathrm{\boldsymbol{\xi}}|=5$ in Eq. 3.3). Figure 4 shows the probability distribution of the trace distance, and Table 4 summarizes the corresponding statistical moments for each model. The distribution of the trace distance shows that the learned UDE models outperform the base model in accuracy over both the training and validation data sets. Here we observe that the affine and the nonlinear neural network models exhibit superior performance when compared to the structure preserving ansatz. This suggests that the Markovian Lindblad model is unable to accurately model all processes present on the QPU, and these processes are better represented by the affine and nonlinear operators. By comparing Tab. 2 and Tab. 4, it is apparent that the learned UDE models capture dynamics on both QPUs to similar accuracy in expectation, but variance on Dev2 is an order of magnitude higher than it is on Dev1. This is attributed to the increased levels of noise present on this QPU. Nevertheless, this demonstrates that our UDE approach is capable of effectively capturing the dynamics of noisy QPUs.
@@
where the magnitude of detuning is similar to those obtained on Dev1. The magnitude of the perturbation for $\sigma_{x}$ and $\sigma_{y}$ is, 1.96 kHz and 10.4 kHz respectively, and are significantly larger than on Dev1. The estimated perturbations to the decoherence times are $\mathcal{T}^{Lind}=\left\{10,8.5\right\}~{}\upmu$s. The perturbed $T_{1}$ decay time is approximately $4.6\upmu$s and is an order of magnitude lower than the decay time estimated using standard characterization protocols. The effective $T_{2}$ dephasing time is approximately $1.6\upmu$s and also lower than what was obtained through standard characterization.
Having demonstrated the UDE approach’s ability to learn generalized operators, we investigate the accuracy gained by tailoring the operator for a particular experiment. Figure 5 displays the expected energy, and probability density of the trace distance for a random experiment; Tab. 5 presents the statistical moments for the different UDE approaches when learning Exp-Spec operators. The expected energy of the density matrices computed by the affine and nonlinear models is more closely aligned with the those obtained from LIE; furthermore, the probability densities are higher and more concentrated for smaller trace distances. Although the accuracy of the Exp-Spec operators over the training data is greater than that of the Exp-Gen operators, the accuracy of the two over the validation set are comparable. The expected energy and densities of the trace distance for additional random experiments can be found in Appendix A; these results demonstrate improved accuracy. This reinforces the conclusion that learning Exp-Spec operators greatly improves accuracy.
@@
where the detuning introduced by the Exp-Spec operator is significantly lower than that estimated by the Exp-Gen operator. The estimated perturbed decoherence times are $\mathcal{T}^{Lind}=\left\{17.9,23.1\right\}~{}\upmu$s, with the perturbed $T_{1}$ decay and $T_{2}$ dephasing times of approximately $7.8\upmu$s and $2.9\upmu$s respectively. The lower $T_{1}$ estimates agree with the fast decaying nature of the expected energy seen in Fig. 5, where for $N=2$, the expected energy is equivalent to the population of the first excited state.
@@ 5 Conclusion
This work demonstrated a data-driven approach for learning latent dynamics of quantum processing units (QPUs). The technique augments a base model of the known dynamics (e.g. Lindblad master equation) with a trained source term that models the unknown dynamics apparent in the training data from QPUs. We presented analysis with two base models: Liouville-von Neumann and Lindblad master equation, and three formulations (ansatze) for this data-driven source term: a structure preserving model, an affine model, and a nonlinear neural network model. The structure preserving ansatz offered interpretability and physically consistent time-evolution of the quantum state, whereas the affine and nonlinear models were neither directly interpretable nor yielded physically consistent evolution. The affine and nonlinear models were made physically consistent by applying a spectral filter with renormalization, a-posteriori. The accuracy and applicability of the different ansatze was investigated using data from two different QPUs at the LLNL’s Quantum Design and Integration Testbed (QuDIT), each with a different level of noise.
We demonstrated that the structure preserving and affine models typically required less training data than the nonlinear model, while achieving improved out-of-distribution accuracy over the base models. Furthermore, the structure preserving model converged to similar operators regardless of the underlying base model used. We also showed that the structure preserving ansatz is more accurate for less-noisy QPUs where the latent dynamics are well described by the Markovian Lindblad equation, whereas the affine model performed well on both QPUs. The nonlinear model was seen to perform well only on more noisy QPUs, and hence, both the affine and nonlinear models are better able to model the noise processes. Despite the lack of interpretability and for general application of learning dynamics and noise processes, the affine model was seen to be more accurate and robust, than the remaining ansatze. Both structure preserving and affine models are applicable for learning dynamics of stable, less-noisy QPUs, and can be used as more accurate numerical models which can serve as efficient emulators of QPUs. The UDE approach identifies latent dynamics and noise processes which can greatly inform hardware development. Furthermore, accurate numerical models of the quantum devices can improve the quality of the control pulses needed for performing quantum operators accurately.
@@ Acknowledgements
This work was supported in part by LLNL Laboratory Directed Research and Development project 23-ERD-038 and was performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under Contract DE-AC52-07NA27344. LLNL-JRNL-858483.
@@ Appendix A Supplementary Material: Learning Dynamics of QPU - Dev2
and the estimated perturbed decoherence times are $\mathcal{T}=\left\{11.2,11.1\right\}~{}\upmu$s with the perturbed $T_{1}$ decay and $T_{2}$ dephasing times of approximately $5.1\upmu$s and $1.9\upmu$s respectively.
@@
and the estimated perturbed decoherence times are $\mathcal{T}=\left\{7.8,6.2\right\}~{}\upmu$s with the perturbed $T_{1}$ decay and $T_{2}$ dephasing times of approximately $3.7\upmu$s and $1.2\upmu$s respectively.